OpenStudy (anonymous):

solve the equation. (1/3)log2^(x+6)=log8^(3x) the answer is 3. show me how and why...

OpenStudy (somethingawesome):

It's not true as written, are you sure you copied the question correctly? If we plug in x = 3, we get 3 log(2) = 27 log(2), which is false.

OpenStudy (anonymous):

well unless i got the answer wrong from my teacher... my answer choices are a) ø b) 3, 0 c) 3 d) 9

OpenStudy (somethingawesome):

Well, maybe I'm misinterpreting it. Do you mean: \[\log(2^{x+6})\] or \[\log_2(x+6)\]

OpenStudy (anonymous):

that second one

OpenStudy (anonymous):

idk how to key in the bases

OpenStudy (somethingawesome):

Ok, that's what my problem was :) it might be clearer to write log_2(x+6). We need to change the base of one of them, and that formula is \[\log_b(x) = \frac{\log_k(x)}{\log_k(b)}\] so \[\log_8(x) = \frac{\log_2(x)}{\log_2(8)} = (1/3) \log_2(x)\] since log_2(8) = 3. Then we now have (1/3) log_2(x+6) = (1/3) log_2(3x) by substituting what we just found. Cancel things out and we have to compare the insides, since they have the same base. x+6 = 3x has the solution x=3.

OpenStudy (anonymous):

thank you... i really appreciate it

OpenStudy (somethingawesome):

Happy to help!